OpenStudy (anonymous):

find (f+g)x, (f-g)x, f(g)x and (t/f)x. f(x)= x^2+3x+2 g(x)=x+1

OpenStudy (amistre64):

f+g; just add at all together; f-g: turn g(x) into -g(x) = -x-1 and combine the two function

OpenStudy (amistre64):

f(g)x ; i think that is saying to multiply the functions together

OpenStudy (amistre64):

x^2 +4x +3 x^2 +2x +1 (x^2 +3x +2)(x+1) = x^3 + 4x^2 +5x +1 f/g can be tricky lol

OpenStudy (amistre64):

g/f is just putting g over f as a fraction and there aint much ulse to do to it; unles it can factor